# Efficient Bayesian Updates for Deep Learning via Laplace Approximations

### v2

@@ 3.2 Fast Approximations of Bayesian Updates for Deep Neural Networks
Our update method is based on a combination of two concepts. First, instead of MC-based BNNs, we suggest using LAs on the last layer of a DNN. Second, we directly modify the approximate posterior distribution of the LA, providing a much more flexible way to adapt it to new data than reweighting. In the following, we explain each component in detail. For now, we focus on binary classification with $K=2$, and refer to Appendix C for an extension to multi-class classification.
@@
$q(\boldsymbol{\omega}|\mathcal{D})=\mathcal{N}(\boldsymbol{\omega}|\hat{{\boldsymbol{\mu}}},\hat{{\boldsymbol{\Sigma}}})\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr\propto\cr\kern 2.0pt\cr\sim\cr\kern-2.0pt\cr}}}q(\boldsymbol{\omega})\prod_{({\boldsymbol{x}},y)\in\mathcal{D}}p(y|{\boldsymbol{x}},\boldsymbol{\omega}),$ (5)
@@
The benefits of using a last-layer LA are manifold. Given access to $q(\boldsymbol{\omega}|\mathcal{D})$ through a Gaussian, we enable more flexible updates by being able to modify the mean and covariance directly. In contrast, MC-based updates only change the approximate distribution by reweighting hypotheses, leading to a strong dependency on the samples $\boldsymbol{\omega}_{1},\dots,\boldsymbol{\omega}_{M}$. Last-layer LAs can be integrated seamlessly into nearly all DNNs, including pretrained models, as only the covariance has to be computed to obtain $q(\boldsymbol{\omega}|\mathcal{D})$. Finally, compared to deep ensembles and MC-Dropout, last-layer LAs introduce minimal computational overhead. While deep ensembles require longer training and MC-dropout impairs the inference time, LAs simply need to calculate a covariance matrix after training and allow fast inference (cf. Eq. (1)) through techniques such as mean-field approximation [32].
Second-Order Update: The second concept focuses on the update step of the Gaussian distribution. When observing new data, we follow the same approach as in Eq. (5), but with $q(\boldsymbol{\omega}|\mathcal{D})$ as our prior:
$q(\boldsymbol{\omega}|\mathcal{D},\mathcal{D}^{\oplus})=\mathcal{N}(\boldsymbol{\omega}|\hat{{\boldsymbol{\mu}}}^{\text{upd}},\hat{{\boldsymbol{\Sigma}}}^{\text{upd}})\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr\propto\cr\kern 2.0pt\cr\sim\cr\kern-2.0pt\cr}}}q(\boldsymbol{\omega}|\mathcal{D})\prod_{({\boldsymbol{x}},y)\in\mathcal{D}^{\oplus}}p(y|{\boldsymbol{x}},\boldsymbol{\omega}),$ (8)
where $\hat{{\boldsymbol{\mu}}}^{\text{upd}}$ and $\hat{{\boldsymbol{\Sigma}}}^{\text{upd}}$ represent the updated mean and covariance, respectively. The resulting updated posterior $q(\boldsymbol{\omega}|\mathcal{D},\mathcal{D}^{\oplus})$ is non-Gaussian due to $p(y|{\boldsymbol{x}},\boldsymbol{\omega})$ being a categorical likelihood. Consequently, the closed-form computation of the integral in Eq. (1) becomes intractable. The basic idea of our update is to approximate the new posterior $q(\boldsymbol{\omega}|\mathcal{D},\mathcal{D}^{\oplus})$ by first applying a second-order optimization step via Gauss-Newton and then estimating the new covariance at that point. Thus, the updated mean and covariance are given by:
@@
where $\overline{\sigma}_{\boldsymbol{x}}={\boldsymbol{h}}_{\boldsymbol{x}}^{\mathrm{T}}{\boldsymbol{\Sigma}}{\boldsymbol{h}}_{\boldsymbol{x}}$ is the predictive variance. The derivation can be found in Appendix D.
The idea behind using second-order optimization techniques is that they are more robust than first-order gradient optimization techniques due to the incorporation of curvature information of the log posterior. This results in a more accurate representation of the loss landscape, leading to improved parameter updates. A critical aspect of our method’s efficiency is that we do not need to recompute the Hessian from scratch. Instead, our updates leverage the covariance available through LAs and use the Woodbury identity [55] for closed-form inversion, significantly reducing computational overhead. Further, a common problem with last-layer LAs is that the Hessian can become a bottleneck when dealing with a large number of classes. To address this, we can approximate the Hessian in Eq. (11) by considering a Gaussian likelihood instead of a a multi-class one, as also done in [30, 12].
@@ 4 Bayesian Updating Experiments
In this section, we evaluate the effectiveness and speed of the proposed update by comparing it against competitors on various benchmark datasets for image and text classification. Our code is publicly available at https://github.com/dhuseljic/dal-toolbox.
@@ 4.1 Experimental Setup
Our experimental design is based on the work of [21]. First, we train a DNN on the training dataset $\mathcal{D}$ (baseline). We then use this baseline DNN to perform a last-layer LA and different Bayesian updates on additional instance-label pairs $\mathcal{D}^{\oplus}$ and compare these results to retraining the DNN on the complete dataset $\mathcal{D}\cup\mathcal{D}^{\oplus}$. To ensure a comprehensive evaluation, we evaluate i) the influence of the step size $\gamma$ on chosen validation datasets, ii) the impact of our update at different learning stages of the DNN, iii) the impact of our update with increasing sizes of new arriving datasets, and iv) the time-efficiency of our update by considering the speed-up factor against competitors. For comparison, we consider MC-based updates by sampling 10k hypotheses from the approximate Gaussian posterior $q(\boldsymbol{\omega}|\mathcal{D})$ and the less complex first-order updates only considering gradients (cf. Appendix A). Since the latter does not use the Hessian, this comparison allows us to assess the benefits of using second-order optimization. Moreover, we exclude retraining solely on $\mathcal{D}^{\oplus}$, as we empirically found that it leads to catastrophic forgetting [19]. For reproducibility, we average all performance metrics across 10 repetitions.
@@
Type Dataset Reference # classes
@@
The datasets $\mathcal{D}$ and $\mathcal{D}^{\oplus}$ are randomly sampled from real-world datasets. We use three image and three text benchmark datasets commonly used in literature [16, 39] with varying complexity reflected through different numbers of classes. Table 1 gives an overview. A detailed summary for each dataset is provided in Appendix E.
The goal of an update method is to ensure both effectiveness and speed. To assess this, we use different performance metrics. For evaluating effectiveness, or how well an update or retraining generalizes, we measure accuracy. When experimenting with hyperparameters, accuracy is assessed on a 10% validation split. Otherwise, it is measured on the test dataset. An optimal update method should achieve the same performance as completely retraining the DNN with $\mathcal{D}\cup\mathcal{D}^{\oplus}$. To assess the speed of an update, we report the speed-up factor compared to retraining by dividing the time required for retraining by the time required for updating. Retraining and updating times were recorded on an NVIDIA RTX 4090 GPU and an AMD Ryzen 9 7950X CPU, respectively.
We choose commonly employed DNN architectures from the literature [16, 50]. For image datasets, we employ a Vision Transformer (ViT) [10] with pretrained weights via self-supervised learning, complemented by a randomly initialized fully connected layer. Specifically, we use the DINOv2-ViT-S/14 model [36] with a feature dimension of $D=384$ in its final hidden layer. For text datasets, we employ the transformer-based pre-trained language model BERT [9]. We utilize bert-based-uncased from the Huggingface library [54] with the [cls] token as the embedding of an instance with a feature dimension of $D=768$ and a maximum sequence length of 512. We train each DNN by finetuning for 200 epochs, employing the Rectified Adam optimizer [31] with a training batch size of 64, a learning rate of 0.01 for images and 0.1 for text, and weight decay of 0.0001. In addition, we utilize a cosine annealing learning rate scheduler. These hyperparameters were determined empirically to be effective across all datasets by investigating the loss convergence on validation splits.
@@ 4.2 Experiments
Hyperparameter Ablation:
Accuracy
Size of $\mathcal{D}^{\oplus}$
Refer to captionSize of $\mathcal{D}^{\oplus}$
@@
In Eq. (9), we introduced the hyperparameter $\gamma$, which controls the step size of our update. Intuitively, this factor determines the extent to which the DNN is influenced by the new dataset $\mathcal{D}^{\oplus}$. This factor is essential to control the update process and avoid issues such as catastrophic forgetting. Similarly, first-order and MC-based updates also utilize this factor to mitigate such problems. For further details, we refer to Appendix A.
@@
To investigate the influence of $\gamma$ and determine a suitable value for all subsequent experiments, we conduct a simple ablation study on two datasets. The results of our update are shown here, while the results for first-order and MC-based updates can be found in Appendix B. We determine the value of $\gamma$ in this manner since an extensive hyperparameter search for update methods is typically impractical in an online setting [8]. Hence, fixing a value beforehand is necessary. We randomly sample an initial dataset $\mathcal{D}$ of 50 instances and train our baseline DNN. Subsequently, updates and retraining are performed on randomly sampled datasets $|\mathcal{D}^{\oplus}|\in\{1,\dots<,10\}$, and the accuracy is computed on a validation split. We repeat this process for different values of $\gamma$.
@@
The resulting curves in Fig. 3 indicate that our update with $\mathcal{D}^{\oplus}$ consistently achieves better performance than the baseline DNN that is only trained on $\mathcal{D}$. For both CIFAR-10 and DBPedia, updating with $\gamma=1$ does not yield accuracies close to retraining, suggesting that the update is too weak. By increasing $\gamma$, we observe accuracies much closer to complete retraining, with $\gamma=10$ being sufficient for CIFAR-10 and DBPedia. For CIFAR-10, we also notice that a very high value, i.e., $\gamma=30$, can lead to worse performance, likely due to catastrophic forgetting. To ensure effective updates across all datasets, we will be using $\gamma=10$ in all subsequent experiments. While this may not be optimal for some datasets, it should ensure a consistently working update process in all cases.
Accuracy Improvement
Size of $\mathcal{D}$
Refer to captionSize of $\mathcal{D}$
Refer to captionSize of $\mathcal{D}$
Accuracy Improvement
Size of $\mathcal{D}$
Refer to captionSize of $\mathcal{D}$
Refer to captionSize of $\mathcal{D}$
@@
Figure 4: Accuracy improvement curves for six benchmark datasets, showing the difference in accuracy between retrained and updated DNNs for varying sizes of $\mathcal{D}$.
@@
Different Learning Stages: To investigate how our update behaves at different stages of learning, we train the baseline DNN on varying sizes of initial datasets $\mathcal{D}$ and update it with a new dataset of fixed size $|\mathcal{D}^{\oplus}|=10$. To better visualize the differences, we report accuracy improvement of updated/retrained DNNs relative to the baseline in Fig. 4. The results demonstrate that our updates provide the highest accuracy improvements across all datasets, highlighting the effective and consistent performance improvements of our update at different stages of learning. While first-order and MC-based updates are also effective in earlier stages (when $|\mathcal{D}|<50$), they tend to be less effective and even deteriorate accuracy in later stages. Compared to the first-order update, our update consistently enhances performance due to including the Hessian. As the Hessian considers curvature information about the posterior, the update is more robust regarding the choice of $\gamma$.
Varying Size of $\mathcal{D}^{\oplus}$: To investigate our update’s behavior with an increasing number of new data points in $\mathcal{D}^{\oplus}$, we train a baseline DNN with a fixed initial dataset $|\mathcal{D}|=100$ and vary the size of the new dataset $|\mathcal{D}^{\oplus}|\in\{10,20,\dots,100\}$. We report the results for the most complex datasets DTD, Banking-77 and Clinc-150. In Fig. 5 we observe that as the size of $\mathcal{D}^{\oplus}$ increases, the accuracy of retraining, our update, and the first-order update consistently improves. In contrast, MC-based updates result in worse accuracies than the baseline, indicating that it is not suited for an increasing size of $\mathcal{D}^{\oplus}$. Considering our update, we see that it consistently achieves better accuracies compared to competitors, regardless of the complexity of the dataset. Moreover, first-order updates seem to be less effective on the more complex datasets such as Banking-77 and Clinc-150, highlighting the importance of the Hessian.
Accuracy
Size of $\mathcal{D}^{\oplus}$
Refer to captionSize of $\mathcal{D}^{\oplus}$
Refer to captionSize of $\mathcal{D}^{\oplus}$
@@
Refer to captionSize of $\mathcal{D}$
Speed-up Factor
Figure 6: Speed-up of update methods compared to retraining.
@@
Time Comparison: Finally, to evaluate the speed of our update, we fix the size of the new dataset to $|\mathcal{D}^{\oplus}|=10$ and compute the speed-up relative to retraining by varying the initial dataset size $|\mathcal{D}|$. Figure 6 presents the speed-up factors of all update methods on CIFAR-10. All methods are faster than retraining, with the first-order update being the fastest. For example, with an initial dataset size of $|\mathcal{D}|=1000$, the first-order update is about 1700 times faster than retraining. Notably, our update provides a similar speed-up factor while yielding more effective updates by using the closed-form update of the Hessian. Compared to MC-based updates, both the first-order and our update are significantly faster.
5 Use Case: Deep Active Learning
To demonstrate the potential of our proposed update in a practical setting, we return to deep AL [48], where a selection strategy iteratively selects unlabeled instances to be labeled by an oracle. This selection is based on information provided by a DNN trained on the currently available labeled instances. The goal is to maximize the DNN’s performance while minimizing the number of label acquisitions. In the deep learning setting, instances are typically chosen in batches to reduce the number of retraining processes after a selection step [20].
The naive idea of using sequential selection strategies for batch selection is to use the top-$b$ instances. As this might cause a lack of diversity between selected instances, batch selection strategies have been proposed to solve this problem [47, 1, 20]. Our idea is to overcome the necessity of batch strategies by using the proposed update method with sequential selection strategies as a fast alternative to retraining. After acquiring $b$ labels, we retrain the DNN similar to batch selection strategies. An algorithm implementing this idea can be found in Appendix F. The hypothesis is that our idea achieves higher performance compared to selecting the top-$b$ instance for already well-performing sequential selection strategies [41], which are Margin [3] and BALD [14]. Additionally, we are interested in whether our idea can act as a replacement for the diversity component of batch selection strategies. Therefore, we also evaluate the popular strategy Badge [1] in combination with our updates.
Accuracy Improvement
Number of Labels
Refer to captionNumber of Labels
Refer to captionNumber of Labels

### v3

@@ 3.2 Fast Approximations of Bayesian Updates for Deep Neural Networks
Our update method is based on a combination of two concepts. First, instead of MC-based BNNs, we suggest using LAs on the last layer of a DNN. Second, we directly modify the approximate posterior distribution of the LA, providing a much more flexible way to adapt it to new data than reweighting. In the following, we explain each component in detail. For now, we focus on binary classification with $K=2$, and refer to Appendix˜C for an extension to multi-class classification.
@@
$q(\boldsymbol{\omega}|\mathcal{D})=\mathcal{N}(\boldsymbol{\omega}|\hat{{\boldsymbol{\mu}}},\hat{{\boldsymbol{\Sigma}}})\mathrel{\vbox{\halign{\hfil$#$\cr\propto\cr\kern 2.0pt\cr\sim\cr\kern-2.0pt\cr}}}q(\boldsymbol{\omega})\prod_{({\boldsymbol{x}},y)\in\mathcal{D}}p(y|{\boldsymbol{x}},\boldsymbol{\omega}),$ (5)
@@
The benefits of using a last-layer LA are manifold. Given access to $q(\boldsymbol{\omega}|\mathcal{D})$ through a Gaussian, we enable more flexible updates compared to MC-based ones, as we can directly modify the mean and covariance. In contrast, MC-based updates only change the approximate distribution by reweighting hypotheses, leading to a strong dependency on the samples $\boldsymbol{\omega}_{1},\dots,\boldsymbol{\omega}_{M}$. Last-layer LAs can be integrated seamlessly into nearly all DNNs, including pretrained models, as only the covariance has to be computed to obtain $q(\boldsymbol{\omega}|\mathcal{D})$. This is particularly important in deep AL, where recent findings highlight self-supervised learning as a crucial factor in selecting informative instances [16, 15]. Finally, compared to deep ensembles and MC-Dropout, last-layer LAs introduce minimal computational overhead. While deep ensembles require longer training and MC-dropout impairs the inference time, LAs simply need to calculate a covariance matrix after training and allow fast inference through techniques such as mean-field approximation [28].
Second-Order Update: The second concept focuses on the update step of the Gaussian distribution. Observing new data, we follow the same approach as in Eq.˜5, but with $q(\boldsymbol{\omega}|\mathcal{D})$ as our prior:
$q(\boldsymbol{\omega}|\mathcal{D},\mathcal{D}^{\oplus})=\mathcal{N}(\boldsymbol{\omega}|\hat{{\boldsymbol{\mu}}}^{\text{upd}},\hat{{\boldsymbol{\Sigma}}}^{\text{upd}})\mathrel{\vbox{\halign{\hfil$#$\cr\propto\cr\kern 2.0pt\cr\sim\cr\kern-2.0pt\cr}}}q(\boldsymbol{\omega}|\mathcal{D})\prod_{({\boldsymbol{x}},y)\in\mathcal{D}^{\oplus}}p(y|{\boldsymbol{x}},\boldsymbol{\omega}),$ (8)
where $\hat{{\boldsymbol{\mu}}}^{\text{upd}}$ and $\hat{{\boldsymbol{\Sigma}}}^{\text{upd}}$ represent the updated mean and covariance, respectively. The resulting updated posterior $q(\boldsymbol{\omega}|\mathcal{D},\mathcal{D}^{\oplus})$ is non-Gaussian due to $p(y|{\boldsymbol{x}},\boldsymbol{\omega})$ being a categorical likelihood. Consequently, the closed-form computation of the integral in Eq.˜1 becomes intractable. The basic idea of our update is to approximate the new posterior $q(\boldsymbol{\omega}|\mathcal{D},\mathcal{D}^{\oplus})$ by first applying a second-order optimization step via Gauss-Newton and then estimating the new covariance at that point. Thus, the updated mean and covariance are given by:
@@
where $\overline{\sigma}_{\boldsymbol{x}}={\boldsymbol{h}}_{\boldsymbol{x}}^{\mathrm{T}}{\boldsymbol{\Sigma}}{\boldsymbol{h}}_{\boldsymbol{x}}$ is the predictive variance. The derivation can be found in Appendix˜D.
The idea behind using second-order optimization techniques is that they are more robust than first-order gradient optimization techniques due to the incorporation of curvature information of the log posterior. This results in a more accurate representation of the loss landscape, enabling more efficient and robust parameter updates that are less sensitive to hyperparameter choices. A critical aspect of our method’s efficiency is that we do not need to recompute the Hessian from scratch. Instead, our updates leverage the covariance available through LAs and use the Woodbury identity [43] for closed-form inversion, significantly reducing computational overhead. Further, a common problem with last-layer LAs is that the Hessian can become a bottleneck when dealing with a large number of classes. To address this, we can approximate the Hessian in Eq.˜11 by considering a Gaussian likelihood instead of a multi-class one, as also done in [26]. Lastly, we want to highlight that an assumption of an LA is that we are at the mode of a distribution, and adding more data violates this assumption. As we focus on AL and only update with a few (up to hundreds) instances at a time, this issue is less severe.
@@ 4 Bayesian Updating Experiments
In this section, we evaluate the efficiency of the proposed update by comparing it against competitors on various benchmark datasets for image and text classification. Our code is publicly available at https://github.com/dhuseljic/dal-toolbox.
@@ 4.1 Experimental Setup
Our experimental design is based on the work of [20]. First, we train a DNN on the training dataset $\mathcal{D}$ (baseline). We then use this baseline DNN to evaluate a last-layer LA and related Bayesian updates on additional instance-label pairs $\mathcal{D}^{\oplus}$ and compare these results to retraining the DNN on the complete dataset $\mathcal{D}\cup\mathcal{D}^{\oplus}$. We evaluate (i) the influence of the step size $\gamma$ on chosen validation datasets, (ii) the impact of our update at different learning stages of the DNN, (iii) the impact of our update with increasing sizes of new arriving datasets, and (iv) the time efficiency of our update by considering the speed-up factor against retraining. For comparison, we consider MC-based updates by sampling 10k hypotheses from the approximate Gaussian posterior $q(\boldsymbol{\omega}|\mathcal{D})$ and the less complex first-order updates only considering gradients. Note that the latter is equivalent to the continual learning strategy of [34], as we demonstrate in Appendix˜A. Since first-order updates do not use the Hessian, this comparison also allows us to assess the benefits of using second-order optimization. We exclude retraining solely on $\mathcal{D}^{\oplus}$, as we empirically found that it leads to catastrophic forgetting [18]. All performance metrics are averaged across 10 repetitions. For visual clarity, we do not report standard errors.
@@
Type Dataset # classes
@@
The datasets $\mathcal{D}$ and $\mathcal{D}^{\oplus}$ are randomly sampled from real-world datasets. We use three image and three text benchmark datasets commonly used in literature [16, 33] with varying complexity reflected through different numbers of classes. Table˜1 gives an overview. A detailed summary for each dataset is provided in Appendix˜E.
The goal of an update method is to ensure both effectiveness and speed. To assess this, we use different performance metrics. To evaluate effectiveness, or how well an update or retraining generalizes, we measure accuracy. When experimenting with hyperparameters, accuracy is assessed on a 10% validation split. Otherwise, it is measured on the test dataset. An optimal update method should achieve the same performance as completely retraining the DNN with $\mathcal{D}\cup\mathcal{D}^{\oplus}$. To assess the speed of an update, we report the speed-up factor compared to retraining by dividing the time required for retraining by the time required for updating (Eqs.˜10 and 11). Retraining and updating times were recorded on an NVIDIA RTX 4090 GPU and an AMD Ryzen 9 7950X CPU, respectively.
We choose common pretrained DNN architectures from the literature [16, 15]. For image datasets, we employ a Vision Transformer (ViT) [11] with pretrained weights via self-supervised learning, complemented by a randomly initialized fully connected layer. Specifically, we use the DINOv2-ViT-S/14 model [31] with a feature dimension of $D=384$ in its final hidden layer. For text datasets, we employ the transformer-based pretrained language model BERT [10]. We utilize bert-based-uncased from the Huggingface library [42] with a feature dimension of $D=768$ and a maximum sequence length of 512. We train each DNN by finetuning for 200 epochs, employing the Rectified Adam optimizer [27] with a training batch size of 64, a learning rate of 0.01 for images and 0.1 for text, and weight decay of 0.0001. In addition, we utilize a cosine annealing learning rate scheduler. These hyperparameters were determined empirically to be effective across all datasets by investigating the loss convergence on validation splits.
@@ 4.2 Experiments
Hyperparameter Ablation: In Eq.˜10, we introduced the hyperparameter $\gamma$, which controls the step size of our update. Intuitively, this factor determines the extent to which the DNN is influenced by the new dataset $\mathcal{D}^{\oplus}$. This factor is essential to control the update process and avoid issues such as catastrophic forgetting. Similarly, first-order and MC-based updates also utilize this factor to mitigate such problems. For further details, we refer to Appendix˜A.
@@
To investigate the influence of $\gamma$ and determine a suitable value for all subsequent experiments, we conduct a simple ablation study on two datasets. The results of our update are shown here, while the results for first-order and MC-based updates can be found in Appendix˜B. We determine the value of $\gamma$ in this manner since an extensive hyperparameter search for update methods is typically impractical in an online setting [9]. Hence, fixing a value beforehand is necessary. We randomly sample an initial dataset $\mathcal{D}$ of 50 instances and train our baseline DNN. Subsequently, updates and retraining are performed on randomly sampled datasets $|\mathcal{D}^{\oplus}|\in\{1,\dots,10\}$, and the accuracy is computed on a validation split. We repeat this process for different values of $\gamma$.
@@
The resulting curves in Fig.˜2 indicate that our update with $\mathcal{D}^{\oplus}$ consistently achieves better performance than the baseline DNN that is only trained on $\mathcal{D}$. For both CIFAR-10 and DBPedia, updating with $\gamma=1$ does not yield accuracies close to retraining, suggesting that the update is too weak. By increasing $\gamma$, we observe accuracies much closer to complete retraining, with $\gamma=10$ being sufficient for CIFAR-10 and DBPedia. For CIFAR-10, we also notice that a very high value, i.e., $\gamma=30$, can lead to worse performance, likely due to catastrophic forgetting. To ensure effective updates across all datasets, we will be using $\gamma=10$ in all subsequent experiments. While this may not be optimal for some datasets, it should ensure a consistently working update in all cases.
@@
Figure 3: Accuracy improvement curves for benchmark datasets, showing the difference in accuracy between retrained and updated DNNs for varying sizes of $\mathcal{D}$.
@@
Different Learning Stages: To investigate how our update behaves at different stages of learning, we train the baseline DNN on varying sizes of initial datasets $\mathcal{D}$ and update it with a new dataset of fixed size $|\mathcal{D}^{\oplus}|=10$. To better visualize the differences, we report accuracy improvement of updated/retrained DNNs relative to the baseline in Fig.˜3. The results demonstrate that our updates provide the highest accuracy improvements across all datasets, highlighting the effective and consistent performance improvements of our update at different learning stages. While first-order and MC-based updates are also effective in earlier stages (when $|\mathcal{D}|<50$), they tend to be less effective and even deteriorate accuracy in later stages. Compared to the first-order update, our update consistently enhances performance due to including the Hessian. As the Hessian considers curvature information about the posterior, the update is more robust regarding the choice of $\gamma$.
@@
Figure 5: Speed-up of updates compared to retraining.
Varying Size of $\mathcal{D}^{\oplus}$: To investigate our update’s behavior with an increasing number of new data points in $\mathcal{D}^{\oplus}$, we train a baseline DNN with a fixed initial dataset $|\mathcal{D}|=100$ and vary the size of the new dataset $|\mathcal{D}^{\oplus}|\in\{10,20,\dots,100\}$. We report the results for the most complex datasets DTD, Banking-77, and Clinc-150. In Fig.˜4, we observe that as the size of $\mathcal{D}^{\oplus}$ increases, the accuracy of retraining, our update, and the first-order update consistently improves. In contrast, MC-based updates result in worse accuracies than the baseline, indicating that it is not suited for an increasing size of $\mathcal{D}^{\oplus}$. Considering our update, we see that it consistently achieves better accuracies compared to competitors, regardless of the complexity of the dataset. Moreover, first-order updates seem to be less effective on the more complex datasets such as Banking-77 and Clinc-150, highlighting the importance of the Hessian.
@@
Time Comparison: Finally, to evaluate the speed of updates, we fix the size of the new dataset to $|\mathcal{D}^{\oplus}|=10$ and compute the speed-up relative to retraining by varying the initial dataset size $|\mathcal{D}|$. Figure˜5 presents the speed-up factors on CIFAR-10. All update methods are faster than retraining, with the first-order update being the fastest. For example, with an initial dataset of $|\mathcal{D}|=1000$, the first-order update is about 1700 times faster than retraining. Notably, our update provides a similar speed-up factor while yielding more effective updates by using the closed-form Hessian update. Compared to MC-based updates, both the first-order and our update are significantly faster.
5 Deep Active Learning
In this section, we examine the proposed update in AL. First, we introduce a new framework that uses our updates to exploit label information during batch construction. Essentially, this approach mimics single instance AL, in which the model is retrained after each label acquisition. Next, we employ our update to approximate an optimal look-ahead strategy. Instead of obtaining future performance of the DNN with expensive retraining, we realize this through our update. Here, we average metrics over 30 repetitions to account for reproducibility challenges in AL [30]. Labeling budgets and acquisition sizes differ based on the complexity of a dataset. A more detailed experimental setup and all learning curves, including ones that report absolute values, are available in Appendix˜F.